# Operatorial formulation of crimo-taxis phenomena in a street

### v1

@@ Abstract
In this paper, the ladder operators typical of quantum mechanics are used to construct an operatorial version of a model dubbed crimo-taxis. In a classical framework, the crimo-taxis model is described by reaction-diffusion partial differential equations describing a population divided into three interacting subgroups (ordinary citizens, drug users/dealers, and law enforcement personnel). By means of a quantum-like approach, these interacting subgroups are modeled using annihilation, creation, and number fermionic operators. Their dynamics is governed by a self-adjoint and time-independent Hamiltonian operator, emboding the interactions among the groups. Furthermore, a recent variant of the standard Heisenberg approach, namely $(\mathcal{H},\rho)$–induced dynamics, is also taken into account with the goal of enriching the dynamics. The results of some numerical simulations in a one–dimensional setting are presented and discussed.
Keywords. Operatorial models; $(\mathcal{H},\rho)$–induced dynamics; Heisenberg dynamics; Fermionic operators; Crimo-taxis phenomena.
@@ 1 Introduction
The dynamics of interacting species or populations is a crucial topic in theoretical biology, ecology, medicine, chemistry and other fields [1, 2].
In this paper, we focus on a special problem connected to urban crime, where different subgroups of a population can interact. In particular, our aim is to model the interaction among different subgroups of a population, i.e., ordinary citizens, drug users/dealers, and law enforcement personnel; as the crime phenomena do not occur uniformly in space and time, the various subgroups, due to their local interaction and diffusion processes, can spread in space at different times.
Traditionally, the dynamics of spatially distributed groups is modeled using reaction-diffusion partial differential equations [1, 2], which can exhibit interesting behaviors like pattern formation [3]. In particular, an original model, named crimo-taxis and proposed by Epstein in [1], involves three subgroups living in a one–dimensional spatial region representing a street. The original model does not possess asymptotically stable equilibria and this prevents to study Turing instability [4]; consequently, in [3], two possible variants have been proposed such that Turing instability is allowed and some interesting patterns emerge. The two models investigated in [3] can be written in compact form as
@@
for the first and second variant, respectively, whereas $u(x,t)$, $v(x,t)$, and $w(x,t)$ denote the densities of ordinary citizens (susceptibles), drug users/dealers (infectives), and law enforcement personnel at street position $x$ at time $t$, respectively; $r$ represents the susceptible reproduction rate, $\beta$ the infection rate, $\gamma$ the police arrest rate, $b$ the police natural decay rate, $\xi$ the police growth rate, whereas the coefficients $D_{ij}$ are the self- and cross-diffusion terms ($i=j$ and $i\neq j$, respectively). All parameters involved in (1) are assumed to be constant and positive (see [1, 3] for details).
@@
Both models admit asymptotically stable homogeneous coexistence equilibria, susceptible of losing their stability due to the self- and cross-diffusive terms; moreover, the patterns exhibited by these two variants show a different spatial distribution of the three subgroups [3].
@@
In this paper, we propose an alternative approach using the mathematical framework of quantum mechanics [5, 6], specifically raising, lowering operators and number representation. In particular, we implement an operatorial model where the actors of the system are represented by annihilation, creation and number fermionic operators whose evolution is ruled by a time–independent Hermitian Hamiltonian $\mathcal{H}$ within the recently introduced approach of $(\mathcal{H},\rho)$–induced dynamics (see [7], and references therein).
@@
The operatorial approach, initially developed in quantum mechanics, has been successfully applied to several macroscopic systems outside physics (see [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], and references therein, where the main motivations suggesting the use of the tools originally developed in a quantum context to describe classical situations have been widely discussed).
@@
In this framework, we consider a system made by the three subgroups of a population in a patchy environment, that compete both locally and nonlocally, each one subject to migration phenomena with different mobilities; the actors of the model are represented by fermionic operators in a Hilbert space. This choice is advantageous for two main reasons: technical, because the Hilbert space is finite-dimensional, ensuring bounded operators, practical, because the mean values of the observables over an initial condition can represent the local densities of subgroups in different cells.
@@
It is worth of being remarked that operatorial models for spatially distributed competing populations, where migration effects are included, have been already proposed [25, 26, 27]. However, our model description includes also non-local competition mechanisms (as in [24]). Moreover, in addition to the standard Heisenberg view to dynamics, we use the extension known as $(\mathcal{H},\rho)$–induced dynamics [7, 23, 28]. The latter approach studies the evolution of the system, as ruled by the Hamiltonian, but with the action, at fixed instants, of some rules whose effect is that of modifying the values of some parameters entering the Hamiltonian (without modifying the structure of the Hamiltonian) as a consequence of the evolution of the system itself. In such a way, the model adjusts itself during the evolution; the rule can be seen as a way to describe the change of the attitudes of the actors during the evolution, avoiding the introduction of a time dependent Hamiltonian or the consideration of other elements into the system such as a reservoir [29], with associated analytical and computational complications. In some applications (see [23, 28, 30]), the effect of the rules was to allow the system to approach asymptotic equilibrium states. However, we observe that this behavior does not necessarily manifest itself in the spatially distributed model investigated in this paper, even if it introduces a kind of irreversibility in the evolution.
@@
The paper is structured as follows. In Section 2, we fix the notation and provide an operatorial formulation of the crimo-taxis model. The dynamics is governed by a time-independent self-adjoint quadratic Hamiltonian operator, which accounts for both local and nonlocal competition and migration effects. Furthermore, the analytical solution is exhibited. In Section 3, the $(\mathcal{H},\rho)$–induced dynamics approach is briefly reviewed, and two suitable rules acting on inertia parameters only, or on inertia and diffusion parameters, providing also a social interpretation, are introduced. In Section 4, in a one–dimensional domain, two scenarios characterized by different initial spatial distributions of the three subgroups have been considered, and some numerical results, both in the case of the standard Heisenberg dynamics and using the $(\mathcal{H},\rho)$–induced dynamics approach, are presented. Finally, Section 5 contains our concluding remarks, as well as some future extensions of the operatorial model presented in this paper.
@@ 2 Operatorial framework
In this Section, we set the notation and construct the operatorial formulation of the crimo-taxis model, starting from a general description of its components. We introduce the fermionic operators $a_{j,\alpha}$ and $a_{j,\alpha}^{\dagger}$ ($j=1,2,3$, $\alpha=1,\ldots,N$) to represent the actors of the model (ordinary citizens, drug users/dealers and law enforcement personnel) in the generic cell $\alpha$ of a one–dimensional region $\Omega=[0,L]$ representing a street of length $L$. We note that the symbol $\dagger$ denote the adjoint operation, while $a_{j,\alpha}$ and $a^{\dagger}_{j,\alpha}$ denote the annihilation and creation operators, respectively. Furthermore, we need to introduce the number operators [8]
$\hat{n}_{j,\alpha}=a_{j,\alpha}^{\dagger}a_{j,\alpha},\qquad j=1,\ldots,3,\qquad\alpha=1,\ldots,N.$ (2)
The choice (that could seem too much restrictive) of considering a one–dimensional spatial region is not just (or only) a trick to simplify the analysis; however, migratory phenomena are often observed along well defined directions. In such circumstances, a model where the three subgroups of population are assumed spatially distributed in a one–dimensional region can result appropriate.
According to the formalism of second quantization, the fermionic operators satisfy the Canonical Anticommutation Relations (CAR):
$\{a_{j,\alpha},a_{k,\beta}\}=0,\qquad\{a^{\dagger}_{j,\alpha},a^{\dagger}_{k,\beta}\}=0,\qquad\{a_{j,\alpha},a^{\dagger}_{k,\beta}\}=\delta_{j,k}\,\delta_{\alpha,\beta}\,\mathcal{I},$ (3)
where $j,k=1,2,3$, $\;\alpha,\beta=1,\ldots,N$, $\;\{A,B\}:=AB+BA$ denotes the anticommutator between the operators $A$ and $B$, $\mathcal{I}$ is the identity operator, and $\delta_{i,j}$ represents the Kronecker symbol.
The Hilbert space $\mathbb{H}$, where the fermionic model representing the system $\mathcal{S}$ lives, is constructed as the linear span of the orthonormal set of vectors
@@
$\varphi_{\mathbf{0},\mathbf{0},\mathbf{0}}\equiv\varphi_{0,0,\ldots,0},$
i.e., the vector $\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}}$ is obtained by acting on the vacuum $\varphi_{\mathbf{0},\mathbf{0},\mathbf{0}}$ with the powers of the operators $a_{j,\alpha}^{\dagger}$, whit $n_{j,\alpha}=0,1$. As a consequence, we get $\hbox{dim}(\mathbb{H})=2^{3N}$.
The vector $\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}}$ means that it is initially assigned a mean value equal to $n_{j,\alpha}$ to the $j$–th population in the cell $\alpha$ such that
@@
The mean values of number operators, say
$n_{j,\alpha}=\langle\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}},\widehat{n}_{j,\alpha}\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}}\rangle,\qquad j=1,\ldots,3,\qquad\alpha=1,\ldots,N$ (7)
are interpreted, once assigned an initial condition, as the local density of the population $\mathcal{P}_{j}$ in the cell $\alpha$.
The above operators are used to construct the Hamiltonian $\mathcal{H}$ associated to the system $\mathcal{S}$, which is the generator of the evolutive dynamics. In details, we will assume that
@@
$\mathcal{H}_{I}=\sum_{\alpha=1}^{N}\lambda_{1,\alpha}(a_{1,\alpha}a_{2,\alpha}^{\dagger}+a_{2,\alpha}a_{1,\alpha}^{\dagger})+\lambda_{2,\alpha}(a_{2,\alpha}a_{3,\alpha}^{\dagger}+a_{3,\alpha}a_{2,\alpha}^{\dagger}),$
@@
$\mathcal{H}_{C}=\sum_{\alpha,\beta=1}^{N}p_{\alpha,\beta}\left(\nu_{1,\alpha}(a_{1,\alpha}a_{2,\beta}^{\dagger}+a_{2,\beta}a_{1,\alpha}^{\dagger})+\nu_{2,\alpha}(a_{2,\alpha}a_{3,\beta}^{\dagger}+a_{3,\beta}a_{2,\alpha}^{\dagger})+\right.$
$\qquad\qquad\left.+\nu_{3,\alpha}(a_{1,\alpha}a_{3,\beta}^{\dagger}+a_{3,\beta}a_{1,\alpha}^{\dagger})\right),$
@@
are the different contributions in $\mathcal{H}$, whose description will be given below. In the expressions (9), the parameters $\omega_{j,\alpha}$, $\lambda_{j,\alpha}$, $\mu_{j,\alpha}$, and $\nu_{j,\alpha}$ ($\alpha=1,\ldots,N$, $j=1,\ldots,3$) are positive constants; moreover, the coefficients $p_{\alpha,\beta}$ ($\alpha,\beta=1,\ldots,N$), symmetric with respect to their indices, are expressed by
@@
$p_{\alpha,\beta}=\begin{cases}1\quad\mbox{if }\beta\;\hbox{denotes a cell in the Moore neighborhood}\;M_{\alpha}\;\hbox{of cell}\;\alpha,\\ 0\quad\hbox{elsewhere}.\end{cases}$
@@
Therefore, the cells in the Moore neighborhood of the cell $\alpha\neq 1,N$ are labeled as $\alpha-1$ and $\alpha+1$, whereas the cells labeled with 1 and $N$ have only one neighbor (labeled $2$ and $N-1$, respectively, see Figure 1).
Figure 1: Moore neighborhood of the cell $\alpha$.
Here, some comments about the various contributions (9) in the Hamiltonian $\mathcal{H}$ are detailed:
  • $\mathcal{H}_{0}$ is the free part of $\mathcal{H}$, and $\omega_{i,\alpha}$ are parameters somehow related to the inertia of the operators associated to the agents of $\mathcal{S}$: in fact, they can be thought of as a measure of the tendency of each degree of freedom to stay constant in time [8].
@@
  • $\mathcal{H}_{I}$ rules the local (i.e., in the same cell $\alpha$) competitive interactions between the three subgroups. In particular, the coefficients $\lambda_{1,\alpha}$ ($\lambda_{2,\alpha}$, respectively) give a measure of the strength of the interactions between $a_{1,\alpha}$ and $a_{2,\alpha}$ ($a_{2,\alpha}$ and $a_{3,\alpha}$, respectively). The contribution $a_{1,\alpha}a_{2,\alpha}^{\dagger}$ is interpreted as a competition term since a particle associated to $a_{1,\alpha}$ is destroyed and a particle associated to $a_{2,\alpha}$ is created; the adjoint term $a_{2,\alpha}a_{1,\alpha}^{\dagger}$ swaps the roles of the two modes.
@@
  • $\mathcal{H}_{D}$ is responsible for the diffusion of the three subgroups in the region; $\mu_{j,\alpha}$ represents the mobility coefficient of the $j$th population, $\mathcal{P}_{j}$, in the cell $\alpha$. For example, the contribution $a_{1,\alpha}a_{1,\beta}^{\dagger}$ is such that for the actor associated to $a_{1,\alpha}$ a particle is destroyed and for the actor associated to $a_{1,\beta}$ a particle is created (in other words, the particle is annihilated in cell $\alpha$ and created in cell $\beta$, where $\beta\in M_{\alpha}$, mimicking the diffusion in a neighboring cell); the adjoint term $a_{1,\beta}a_{1,\alpha}^{\dagger}$ swaps the roles of the two modes.
  • $\mathcal{H}_{C}$ takes into account a nonlocal competitive interaction of the three subgroups, and $\nu_{j,\alpha}$ measures the strength of nonlocal interaction (absent if $\nu_{j,\alpha}=0$); actually, the competition between the three populations occurs in adjacent cells.
We note also that $\mathcal{H}_{I}$ and $\mathcal{H}_{C}$ describe the effects of competition and cross-diffusion typically found in classical reaction-diffusion models [1].
@@
Thence, adopting the Heisenberg representation, the operators $a_{j,\alpha}$ evolve in time according to the following ordinary differential equations:
$\dot{a}_{j,\alpha}=\textrm{i}[\mathcal{H},a_{j,\alpha}],\qquad j=1,\ldots,3,$ (10)
@@
where $[A,B]=AB-BA$ denotes the the commutator between the operators $A$ and $B$. Since the Hamiltonian operator $\mathcal{H}$ is quadratic, it follows that the ordinary differential equations (10) turn out to be linear:
@@
$+\left.\sum_{\beta=1}^{N}p_{\alpha\beta}\left((\mu_{2,\alpha}+\mu_{2,\beta})a_{2,\beta}+\nu_{1,\beta}a_{1,\beta}+\nu_{2,\alpha}a_{3,\beta}\right)\right),$
@@
$+\left.\sum_{\beta=1}^{N}p_{\alpha\beta}((\mu_{3,\alpha}+\mu_{3,\beta})a_{3,\beta}+\nu_{2,\beta}a_{2,\beta}+\nu_{3,\beta}a_{1,\beta})\right).$
In the following Remark, we underline a possible analogy between reaction–diffusion equations and operatorial models.

### v2

@@ Abstract
In this paper, the ladder operators and a quantum-like approach are used to construct an operatorial version of a model dubbed crimo-taxis. In a classical framework, the crimo-taxis model is represented by reaction-diffusion partial differential equations describing a population divided into three interacting subgroups (ordinary citizens, drug users/dealers, and law enforcement personnel). In this new framework, the agents of the subgroups are modeled using annihilation, creation, and number fermionic operators, and their time evolution is assumed to be ruled by a Hermitian time-independent Hamiltonian operator suitable to capture the interactions among the subgroups. Furthermore, a recent extension of the standard Heisenberg view, namely $(\mathcal{H},\rho)$–induced dynamics, is also taken into account. Two scenarios, characterized by different initial spatial distributions of the subgroups, are considered. The results of some numerical simulations in a one–dimensional setting are presented and discussed.
Keywords. Operatorial models; Fermionic operators; Crimo-taxis phenomena; Heisenberg dynamics; $(\mathcal{H},\rho)$–induced dynamics.
@@ 1 Introduction
The collective dynamics like the one where interacting species or populations occupy a given spatial region is a crucial topic in theoretical biology, ecology, social science, …[1, 2].
In this paper, we focus on a special problem connected to urban crime, where different subgroups of a population are interacting. In particular, our aim is to model the resulting dynamics where different subgroups of a population, i.e., ordinary citizens, drug users/dealers, and law enforcement personnel occupy a spatial region like a street; as the crime phenomena do not occur uniformly in space and time, the various subgroups, due to their local interaction and diffusion processes, can spread in space at different times.
Traditionally, the dynamics of spatially distributed groups is modeled using reaction-diffusion partial differential equations [1, 2], which can exhibit complex outcomes like pattern formation [3]. In particular, a model, named crimo-taxis, proposed by Epstein in [1], involves three subgroups living in a one–dimensional spatial region representing a street. In [3], it has been shown that Epstein’s model does not admit asymptotically stable equilibria, preventing one to study Turing instability [4]; consequently, two possible variants have been proposed such that Turing instability is allowed and some interesting patterns have been exhibited. The two models investigated in [3] can be written in compact form as
@@
whereas $u(t,x)$, $v(t,x)$, and $w(t,x)$ denote the densities of ordinary citizens (susceptibles), drug users/dealers (infectives), and law enforcement personnel at the time $t$ in the street position $x$, respectively; $r$ represents the susceptible reproduction rate, $\beta$ the infection rate, $\gamma$ the police arrest rate, $b$ the police natural decay rate, $\xi$ the police growth rate, whereas the coefficients $D_{ij}$ are the self-diffusion and cross-diffusion terms ($i=j$ and $i\neq j$, respectively). All parameters involved in (1) are assumed to be constant and positive (see [1, 3] for details).
@@
The models (1) admit asymptotically stable homogeneous coexistence equilibria, with the possibility of losing their stability due to the self-diffusion and cross-diffusion terms; moreover, the patterns exhibited by these two variants show asymptotically a stable different spatial distribution of the three subgroups [3].
@@
In this paper, we propose an alternative approach using the mathematical framework of quantum mechanics [5, 6], specifically raising, lowering operators and number representation. In particular, we implement an operatorial model where the agents of the system are represented by annihilation, creation and number fermionic operators whose evolution is governed by a Hermitian time–independent Hamiltonian $\mathcal{H}$.
@@
The operatorial approach, initially developed in quantum mechanics, has been successfully applied to several contexts outside physics (see [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], and references therein, where the choice of using quantum-like tools to discuss classical situations has been largely motivated).
@@
In this framework, we consider a system made by the three subgroups of a population living in a street, that can compete both locally and nonlocally (in the same and adiacent cells, respectively), and can spread with different diffusions; we use fermionic operators, living in a Hilbert space, to represent the agents of the model. This choice is advantageous for two main reasons: technical, because the Hilbert space is finite-dimensional, ensuring bounded operators, and practical, because, assigning an initial condition, the mean values of the number operators can represent local densities of subgroups in different cells.
@@
We also notice that the operatorial approaches for spatially distributed competing populations, where migration effects are included, have already been introduced [24, 25, 26]. However, our model description includes also nonlocal competition effects (as in [23]). Furthermore, besides the classical Heisenberg dynamics, we use the extension known as $(\mathcal{H},\rho)$–induced dynamics (see [27, 22, 28], and references therein); this method allows to study the system dynamics, ruled by the Hamiltonian, superposing periodically some rules able to modify the values assumed by some parameters involved in the Hamiltonian (without changing its structure) according to the evolution of the system itself, thus allowing the model to adjust itself during the evolution. The rule can be seen as a way to describe the change of the attitudes of the agents during the evolution, avoiding the use of an explicitly time dependent Hamiltonian operator or the consideration of other elements into the system such as a reservoir [29], with associated analytical and computational complications. In several applications (see [22, 28, 30]), the system approaches asymptotic equilibrium states, due to the effect of the rules. However, it can be observed that this dynamics does not necessarily manifest itself in the model investigated in this paper, even if, by using some rules, a kind of irreversibility in the dynamics is introduced.
@@
The content of the paper is structured as follows. In Section 2, the notation is fixed, and an operatorial formulation of the crimo-taxis model is provided. A time-independent self-adjoint quadratic Hamiltonian operator rules the dynamics, accounting for both local and nonlocal diffusion and competition effects. Furthermore, the analytical solution is exhibited. In Section 3, the extension $(\mathcal{H},\rho)$–induced dynamics is shortly reviewed, and two suitable rules acting on inertia parameters only, and on both inertia and diffusion parameters (susceptible of having a social interpretation) are introduced. In Section 4, in a one–dimensional domain, two scenarios characterized by different initial spatial distributions of the three subgroups have been considered, and some numerical results, both in the case of the standard Heisenberg view and using the $(\mathcal{H},\rho)$–induced dynamics approach, are presented. Finally, Section 5 summarizes our conclusions, as well as further developments of the operatorial model here investigated.
@@ 2 Operatorial framework
In this Section, we set the notation and construct the operatorial formulation of the crimo-taxis model, starting from a general description of its components. We introduce the fermionic operators $a_{j,\alpha}$ and $a_{j,\alpha}^{\dagger}$ ($j=1,2,3$, $\alpha=1,\ldots,N$) to represent the agents of the system (ordinary citizens, drug users/dealers and law enforcement personnel) in the generic cell $\alpha$ of a one–dimensional domain $\Omega$ representing a street of length $L$. Choosing a one–dimensional setting may seem excessively restrictive, or just only an expedient to reduce the amount of computation; however, this assumption is motivated by the fact that, in natural as well social phenomena, diffusion phenomena can be often seen along well defined directions. As a consequence, considering as a spatial domain a one–dimensional street, where the three subgroups of a population are distributed, can be realistic. The three subgroups are denoted with $\mathcal{P}_{j}$ ($j=1,2,3$): $\mathcal{P}_{1}$ refers to ordinary citizens, $\mathcal{P}_{2}$ to drug users/dealers, and $\mathcal{P}_{3}$ to law enforcement personnel. We note also that the symbol indicates the adjoint operation, whereas $a_{j,\alpha}$ and $a^{\dagger}_{j,\alpha}$ are the annihilation and creation operators, respectively. Furthermore, we introduce the self-adjoint number operators (the observables of our system) [7]
$\widehat{n}_{j,\alpha}=a_{j,\alpha}^{\dagger}a_{j,\alpha},\qquad j=1,2,3,\qquad\alpha=1,\ldots,N.$ (2)
For fermionic annihilation and creation operators, the Canonical Anticommutation Relations (CAR) hold true, i.e.,
$\{a_{i,\alpha},a_{j,\beta}\}=0,\qquad\{a^{\dagger}_{i,\alpha},a^{\dagger}_{j,\beta}\}=0,\qquad\{a_{i,\alpha},a^{\dagger}_{j,\beta}\}=\delta_{i,j}\,\delta_{\alpha,\beta}\,\mathcal{I},$ (3)
where $i,j=1,2,3$, $\;\alpha,\beta=1,\ldots,N$, $\;\{A,B\}:=AB+BA$ denotes the anticommutator between the operators $A$ and $B$, $\mathcal{I}$ is the identity operator, and $\delta_{i,j}$ represents the Kronecker symbol.
The ladder operators live in a Hilbert space $\mathbb{H}$ that is represented by the linear span of the orthonormal family of vectors
@@
$\varphi_{\mathbf{0},\mathbf{0},\mathbf{0}}\equiv\varphi_{0,\ldots,0},$
i.e., the vector $\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}}$ is obtained by acting on the vacuum $\varphi_{\mathbf{0},\mathbf{0},\mathbf{0}}$ with the powers of the operators $a_{j,\alpha}^{\dagger}$, being $n_{j,\alpha}=0,1$ for $\alpha=1,\ldots,N$. Then, we get $\hbox{dim}(\mathbb{H})=2^{3N}$.
The vector $\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}}$ means that it is initially assigned a mean value equal to $n_{j,\alpha}$ to the subgroup $\mathcal{P}_{j}$ such that
@@
Finally, the mean values of the number operators, i.e.,
$n_{j,\alpha}=\langle\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}},\widehat{n}_{j,\alpha}\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}}\rangle,\qquad j=1,2,3,\qquad\alpha=1,\ldots,N$ (7)
can be thought, once assigned an initial condition, as the local density of the subgroup $\mathcal{P}_{j}$.
The above operators are used to construct the Hamiltonian $\mathcal{H}$ associated to the system, which is the generator of the evolutive dynamics. In details, we assume that
@@
$\mathcal{H}_{I}$ $=\sum_{\alpha=1}^{N}\left(\lambda_{1,\alpha}(a_{1,\alpha}a_{2,\alpha}^{\dagger}+a_{2,\alpha}a_{1,\alpha}^{\dagger})+\lambda_{2,\alpha}(a_{2,\alpha}a_{3,\alpha}^{\dagger}+a_{3,\alpha}a_{2,\alpha}^{\dagger})\right),$
@@
$\mathcal{H}_{C}$ $=\sum_{\alpha,\beta=1}^{N}p_{\alpha,\beta}\left(\nu_{1,\alpha}(a_{1,\alpha}a_{2,\beta}^{\dagger}+a_{2,\beta}a_{1,\alpha}^{\dagger})+\nu_{2,\alpha}(a_{2,\alpha}a_{3,\beta}^{\dagger}+a_{3,\beta}a_{2,\alpha}^{\dagger})\right.$
$+\left.\nu_{3,\alpha}(a_{1,\alpha}a_{3,\beta}^{\dagger}+a_{3,\beta}a_{1,\alpha}^{\dagger})\right)$
@@
are the different contributions in $\mathcal{H}$, whose description will be given below. In the expressions (9), the parameters $\omega_{j,\alpha}$, $\lambda_{j,\alpha}$, $\mu_{j,\alpha}$ and $\nu_{j,\alpha}$ ($j=1,\ldots,3$, $\alpha=1,\ldots,N$) are costant and positive, whereas the parameters $p_{\alpha,\beta}$ ($\alpha,\beta=1,\ldots,N$), being symmetric with respect to their indices, are expressed by
@@
$p_{\alpha,\beta}=\begin{cases}1\quad\mbox{if }\beta\;\hbox{is adjacent to}\;\alpha,\\ 0\quad\hbox{elsewhere},\end{cases}$
@@
i.e., the terms $p_{\alpha,\beta}$ are non-zero only in a Moore neighborhood. In Figure 1, the one–dimensional Moore neighborhood is shown: the cells at the extremes, i.e., $\alpha=1,N$, are adjacent to only one cell (the cells $\alpha=2$ and $\alpha=N-1$, respectively), and the cells $\alpha=2,\ldots,N-1$ have two adjacent cells (denoted with $\alpha-1$ and $\alpha+1$).
Figure 1: Moore neighborhood in a one–dimensional street.
Below, we detail the quantities (9) entering the Hamiltonian $\mathcal{H}$:
  • the term $\mathcal{H}_{0}$ is the free part of $\mathcal{H}$, and the parameters $\omega_{i,\alpha}$ therein involved are somehow related to the inertia of the corresponding agent [7]; higher values of $\omega_{i,\alpha}$ correspond to a small propensity of the associated agent to not change in time, and vice versa;
@@
  • the term $\mathcal{H}_{I}$ rules the competitive interactions, occurring in the same cell, among the three subgroups. As there is no local competitive interaction between the agents of the subgroups $\mathcal{P}_{1}$ and $\mathcal{P}_{3}$, we use the parameters $\lambda_{1,\alpha}$ ($\lambda_{2,\alpha}$, respectively) to measure the interaction rates between the operators $a_{1,\alpha}$ and $a_{2,\alpha}$ ($a_{2,\alpha}$ and $a_{3,\alpha}$, respectively). Moreover, also the term $a_{1,\alpha}a_{2,\alpha}^{\dagger}$ ($a_{2,\alpha}a_{3,\alpha}^{\dagger}$, respectively) can be interpreted as a competitive one since a particle associated to $a_{1,\alpha}$ ($a_{2,\alpha}$, respectively) is destroyed and a particle associated to $a_{2,\alpha}$ ($a_{3,\alpha}$, respectively) is created; the adjoint terms $a_{2,\alpha}a_{1,\alpha}^{\dagger}$ and $a_{3,\alpha}a_{2,\alpha}^{\dagger}$ swap the roles of the two agents;
@@
  • the term $\mathcal{H}_{D}$ accounts for the diffusion of the three subgroups in the spatial domain; the parameters $\mu_{j,\alpha}$ represent the diffusion coefficients of the $j$th subgroup in the cell $\alpha$. Also, for each agent, the terms $a_{j,\alpha}a_{j,\beta}^{\dagger}$ mean that a particle associated to $a_{j,\alpha}$ is destroyed and a particle associated to $a_{j,\beta}$ is created (in other words, the particle is annihilated in the cell $\alpha$ and created in the cell $\beta$, mimicking the diffusion in a neighboring cell); moreover, the adjoint terms $a_{j,\beta}a_{j,\alpha}^{\dagger}$ exchange the agents role;
  • the term $\mathcal{H}_{C}$ represents the nonlocal competitive interactions among the three subgroups, taking into account that their competition takes place only in adjacent cells; the values assumed by the parameters $\nu_{j,\alpha}$ allow us to measure the strength of these interactions.
@@
Thence, according to the Heisenberg view, the time evolution of the annihilation operators $a_{j,\alpha}$ is ruled by the following ordinary differential equations:
$\dot{a}_{j,\alpha}=\textrm{i}[\mathcal{H},a_{j,\alpha}],\qquad j=1,2,3,$ (10)
@@
where $[A,B]=AB-BA$ denotes the commutator between the operators $A$ and $B$. Since the Hamiltonian operator $\mathcal{H}$ is quadratic, it follows that the ordinary differential equations (10) turn out to be linear:
@@
$\left.+\sum_{\beta=1}^{N}p_{\alpha\beta}\left((\mu_{2,\alpha}+\mu_{2,\beta})a_{2,\beta}+\nu_{1,\beta}a_{1,\beta}+\nu_{2,\alpha}a_{3,\beta}\right)\right),$
@@
$\left.+\sum_{\beta=1}^{N}p_{\alpha\beta}((\mu_{3,\alpha}+\mu_{3,\beta})a_{3,\beta}+\nu_{2,\beta}a_{2,\beta}+\nu_{3,\beta}a_{1,\beta})\right).$
We note also that $\mathcal{H}_{I}$ and $\mathcal{H}_{C}$ describe the effects of competition and cross-diffusion typically found in classical reaction-diffusion models [1]. In the following Remark, we underline a possible analogy between reaction–diffusion equations and operatorial models.